# Vulnerability of Text-Matching in ML/AI Conference Reviewer Assignments to Collusions

### v1

@@ A.1 The IncludeThemes Algorithm
Input: paper title $t_{p}$, paper abstract $a_{p}$, and adversarial reviewer archive $\mathcal{Q}_{r}^{\text{adv}}$
      Output: $a^{\text{adv}}_{p}$, an adversarial abstract that increases similarity to reviewer $r$ by adding themes from $\mathcal{Q}_{r}^{\text{adv}}$.
1:function IncludeThemes($t_{p}$, $a_{p}$, $\mathcal{Q}_{r}^{\text{adv}}$)
@@
4:         if Early stopping is used and EarlyStoppingCheck(${a^{\text{adv}}_{p}}^{(i-1)}$,$\mathcal{Q}_{r}^{\text{adv}}$)  then
@@
8:              $a^{\text{adv}}_{p}\leftarrow$ A modified abstract including the main themes from $\mathcal{Q}_{r}^{\text{adv}}$ into the original abstract $a_{p}$. The human author can modify the abstract to add themes from $\mathcal{Q}_{r}^{\text{adv}}$ or add manual edits to the an LLM-generated version (see prompt in Appendix B.1.1).
9:              $\triangleright$ Make incremental edits to $a^{\text{adv}}_{p}$.
@@
11:                  if SimilarityCheck($t_{p}$,${a^{\text{adv}}_{p}}^{\prime}$,$a^{\text{adv}}_{p}$,$\mathcal{Q}_{r}^{\text{adv}}$,$\delta$) is true  then
12:                       $a^{\text{adv}}_{p}\leftarrow{a^{\text{adv}}_{p}}^{\prime}$triangleright Update $a^{\text{adv}}_{p}$ if the similarity after edits do not drop dramatically.
@@
15:         $\triangleright$ fully automatic mode
16:              $a^{\text{adv}}_{p}\leftarrow$ An LLM generated adversarial abstract that includes themes from $\mathcal{Q}_{r}^{\text{adv}}$ into the original abstract $a_{p}$. To reduce nonsensical abstract generations, we prompt the LLM to follow a format – the generated abstract should say it takes inspiration from ideas in $\mathcal{Q}_{r}^{\text{adv}}$ in one sentence (Appendix B.1.2).
@@
20:     return ${a^{\text{adv}}_{p}}^{(j)}\in\operatorname*{arg\,max}_{i\in[\bm{N}]}s((t_{p},{a^{\text{adv}}_{p}}^{(i)}),\mathcal{Q}_{r}^{\text{adv}})$ of all $i=0,\ldots,\bm{N}$
@@
In this section, we present the formalized algorithm for the IncludeThemes operation. Both the human-in-the-loop and fully automatic modes are documented in Algorithm 1. In the human-in-the-loop mode, we have a hyperparameter $\delta$, which is a small positive value that allows ${a^{\text{adv}}_{p}}^{\prime}$ to have slightly lower similarity than the $a^{\text{adv}}_{p}$ in return for ensuring coherent and consistent adversarial abstracts. We subjectively pick $\delta$ to be a small value (around $0.01$) without systematic tuning.
@@ A.2 The InsertKeywords Algorithm
In this section, we present the formalized algorithm for the InsertKeywords operation. Both the human-in-the-loop and fully automatic modes are documented in Algorithm 2. In addition, we present the FindKeywords subroutine.
@@
Input: $t_{p}$ (the title of $p$), $a^{\text{adv}}_{p}$ (the adversarial abstract from IncludeThemes), and $\mathcal{Q}_{r}^{\text{adv}}$ (the adversarial archive of reviewer $r$)
      Output: $a^{\text{adv}}_{p}$, an adversarial abstract that increases similarity to $r$ by adding keywords from $\mathcal{Q}_{r}^{\text{adv}}$.
1:function InsertKeywords($t_{p}$, $a^{\text{adv}}_{p}$, $\mathcal{Q}_{r}^{\text{adv}}$)
@@
4:         if Early stopping is used and EarlyStoppingCheck(${a^{\text{adv}}_{p}}^{(i-1)}$,$\mathcal{Q}_{r}^{\text{adv}}$then
@@
7:         keywords $\leftarrow$ FindKeywords($t_{p}$, $a^{\text{adv}}_{p}$, $\mathcal{Q}_{r}^{\text{adv}}$, ${\bm{K}}$)
@@
12:                       if ConstraintCheck(${a^{\text{adv}}_{p}}^{\prime}$) is true and SimilarityCheck($t_{p}$,${a^{\text{adv}}_{p}}^{\prime}$,$a^{\text{adv}}_{p}$,$\mathcal{Q}_{r}^{\text{adv}}$,00) is true then
@@
17:         $\triangleright$ fully automatic mode.
18:              $a^{\text{adv}}_{p}\leftarrow$ Adversarial abstract generated by LLM to incorporate all keywords into the current $a^{\text{adv}}_{p}$. The LLM is prompted to leave out any keywords it considers too technical and unrelated to the main topics in $a^{\text{adv}}_{p}$ (the prompt can be found in Appendix B.2).
@@
22:     return ${a^{\text{adv}}_{p}}^{(i)}$ that has highest $s((t_{p},{a^{\text{adv}}_{p}}^{(i)}),\mathcal{Q}_{r}^{\text{adv}})$
@@ A.2.1 FindKeywords Subroutine
We propose the FindKeywords subroutine, which is a greedy search to find keywords that when inserted into $a^{\text{adv}}_{p}$ raise the similarity to the $\mathcal{Q}_{r}^{\text{adv}}$. In order to make this search efficient, we narrow down keywords by using a heuristic that measures the increase in similarity upon appending the word to the current $a^{\text{adv}}_{p}$. Algorithm 3 details the precise algorithm. This is just one simple instantiation of a possible attack strategy that uses the openly available SPECTER weights—more sophisticated attacks are possible, but we find that our simple heuristics are already extremely successful at breaking current systems.
@@
Input: Paper title $t_{p}$, adversarial abstract from IncludeThemes $a^{\text{adv}}_{p}$, adversarial archive $\mathcal{Q}_{r}^{\text{adv}}$, ${\bm{K}}$ (number of keywords to return), and optionally Filter($\cdot$) (a function to filter out undesirable keywords)
@@
1:function FindKeywords($t_{p},a^{\text{adv}}_{p},\mathcal{Q}_{r}^{\text{adv}},{\bm{K}}$)
2:     $W\leftarrow$ All words in titles and abstracts from $\mathcal{Q}_{r}^{\text{adv}}$.
@@
5:     keywordsSimilarity $\leftarrow$ $s((t_{p},a^{\text{adv}}_{p}),\mathcal{Q}_{r}^{\text{adv}})$triangleright Keeps track of estimated similarity to $r$
@@
8:               ${a^{\text{adv}}_{p}}^{(j)}\leftarrow$ adversarial abstract with words appended at the end “{$a^{\text{adv}}_{p}{}$} {$w_{\max}^{(0)}$} … {$w_{\max}^{(i-1)}$} {$w_{j}$}”.
@@
10:         Let $w_{\max}^{(i)}$ be a word $w_{j}$ such that the associated $s((t_{p},{a^{\text{adv}}_{p}}^{(j)}),\mathcal{Q}_{r}^{\text{adv}})$ is the highest in $W$.
[…]
@@
Appendix B LLM Prompts
B.1 IncludeThemes Prompts
@@
B.1.1 Human-in-the-loop Mode
@@ B.1.2 Automatic Mode
For both human-in-the-loop and manual modes, we provide two examples to the LLM to guide its generation. The examples can be found on GitHub (https://github.com/passionfruit03/reviewer_assignments_vulnerability).
B.2 The InsertKeywords Prompt
@@
For this operation, we also provide two examples to the LLM to guide its generation. The examples can be found on GitHub (https://github.com/passionfruit03/reviewer_assignments_vulnerability).
B.3 Prompts used in Section 5.8 (Human Subject Experiment)
As mentioned in Section 5.8, the prompts used for automatic abstract modification in the human subject experiment is tuned separately on the NeurIPS 2023 test data. Although in different words, the IncludeThemes prompt here has the same rules as the automatic mode prompt presented in Appendix B.1.2 about the number of sentences to add, the inspired by format to follow, and using a scientific writing style. Similarly, for the InsertKeywords prompt, the rules here and in Appendix B.2 about how to insert the keywords, use professional writing styles, and not to add completely unrelated keywords are the same. IncludeThemes Prompt (human subject experiment, fully automatic mode): Please help me edit my abstract’s introduction to explain in one sentence that my work is inspired by the list of previous works in the provided JSON dictionary under the key ”related previous works”. Do not include the titles of previous works. Use the usual writing style of technical academic abstracts, avoid exaggerations and figurative language. Do not use flowery words or phrases such as “prowess”. In addition, please remove any identifiable information (e.g. GitHub URLs) in my abstract by simply replacing them with [omitted for de-identification]. I will provide you with a JSON dictionary with the following structure: { ”title”: my paper’s title, ”abstract”: my paper’s abstract ”related previous works”: [ ”title”: title1, ”abstract”: abstract1, …. ] } Format your answer into JSON with the following schema: { ”title”: title string (should be the same as original), ”abstract”: edited abstract with short introduction, explaining my work is inspired by the previous works provided }

### v2

@@ A.1 The IncludeThemes Algorithm
Algorithm 1 formalizes the IncludeThemes operation, and it shows both human-in-the-loop and fully automatic modes. In human-in-the-loop mode, authors can incrementally edit abstracts until coherent and consistent while maintianing high similarity. We have $\delta$, which is a small positive value that allows such incrementally edited $a^{\text{adv}}_{p}$$^{\prime}$ to trade slightly lower similarity than the current $a^{\text{adv}}_{p}$ for ensuring unsuspicious adversarial abstracts. We subjectively pick $\delta$ to be a small value (around $0.01$) without systematic tuning.
@@
Input: paper title $t_{p}$, paper abstract $a_{p}$, and adversarial reviewer archive $Q_{r}^{\text{adv}}$
   Output: $a^{\text{adv}}_{p}$, an adversarial abstract that increases similarity to reviewer $r$ by adding themes from $Q_{r}^{\text{adv}}$.
1:function IncludeThemes($t_{p}$, $a_{p}$, $Q_{r}^{\text{adv}}$)
@@
[…]
@@
8:     $a^{\text{adv}}_{p}\leftarrow$ A modified abstract including the main themes from $Q_{r}^{\text{adv}}$ into the original abstract $a_{p}$. The human author can modify the abstract to add themes from $Q_{r}^{\text{adv}}$ or add manual edits to the an LLM-generated version (see prompt in Appendix E.1.1).
[…]
@@
11:      if SimilarityCheck($t_{p}$$a^{\text{adv}}_{p}$$^{\prime}$$a^{\text{adv}}_{p}$$Q_{r}^{\text{adv}}$$\delta$) is true  then
12:        $a^{\text{adv}}_{p}\leftarrow{a^{\text{adv}}_{p}}^{\prime}$ $\triangleright$ Update $a^{\text{adv}}_{p}$ if the similarity after edits do not drop dramatically.
@@
[…]
16:     $a^{\text{adv}}_{p}\leftarrow$ An LLM generated adversarial abstract that includes themes from $Q_{r}^{\text{adv}}$ into the original abstract $a_{p}$. To reduce nonsensical abstract generations, we prompt the LLM to follow a format – the generated abstract should say it takes inspiration from ideas in $Q_{r}^{\text{adv}}$ in one sentence (Appendix E.1.2).
@@
20:  return ${a^{\text{adv}}_{p}}^{(j)}\in\operatorname*{arg\,max}_{i\in[\bm{N}]}s((t_{p},{a^{\text{adv}}_{p}}^{(i)}),Q_{r}^{\text{adv}})$ of all $i=0,\ldots,\bm{N}$
@@ A.2 The InsertKeywords Algorithm
Algorithm 2 formalizes the InsertKeywords operation, and it also shows both human-in-the-loop and fully automatic modes. We present FindKeywords subroutine separately in the next part (Appendix A.2.1).
@@
Input: $t_{p}$ (the title of $p$), $a^{\text{adv}}_{p}$ (the adversarial abstract from IncludeThemes), and $Q_{r}^{\text{adv}}$ (the adversarial archive of reviewer $r$)
   Output: $a^{\text{adv}}_{p}$, an adversarial abstract that increases similarity to $r$ by adding keywords from $Q_{r}^{\text{adv}}$.
1:function InsertKeywords($t_{p}$, $a^{\text{adv}}_{p}$, $Q_{r}^{\text{adv}}$)
@@
[…]
@@
7:   keywords $\leftarrow$ FindKeywords($t_{p}$, $a^{\text{adv}}_{p}$, $Q_{r}^{\text{adv}}$, $\bm{K}$}) $\triangleright$ Algorithm 3
@@
12:        if ConstraintCheck($a^{\text{adv}}_{p}$$^{\prime}$) is true and SimilarityCheck($t_{p}$$a^{\text{adv}}_{p}$$^{\prime}$$a^{\text{adv}}_{p}$$Q_{r}^{\text{adv}}$,0) is true then
@@
[…]
18:     $a^{\text{adv}}_{p}\leftarrow$ Adversarial abstract generated by LLM to incorporate all keywords into the current $a^{\text{adv}}_{p}$. The LLM is prompted to leave out any keywords it considers too technical and unrelated to the main topics in $a^{\text{adv}}_{p}$ (the prompt can be found in Appendix E.2).
@@
22:  return ${a^{\text{adv}}_{p}}^{(i)}$ that has highest $s((t_{p},{a^{\text{adv}}_{p}}^{(i)}),Q_{r}^{\text{adv}})$
@@ A.2.1 FindKeywords Subroutine
In Algorithm 3, we propose the FindKeywords subroutine, which is a greedy search to find keywords that when inserted into $a^{\text{adv}}_{p}$ raise the similarity to the $Q_{r}^{\text{adv}}$. To make this search efficient, we narrow down keywords by a heuristic that measures the increase in similarity upon appending the word to the current $a^{\text{adv}}_{p}$. This is just one simple instantiation of a possible attack strategy that uses the openly available SPECTER weights—more sophisticated attacks are possible, but we find that our simple heuristics are already extremely successful at breaking current systems.
@@
Input: Paper title $t_{p}$, adversarial abstract from IncludeThemes $a^{\text{adv}}_{p}$, adversarial archive $Q_{r}^{\text{adv}}$, $\bm{K}$} (number of keywords to return), and optionally Filter($\cdot$) (a function to filter out undesirable keywords)
@@
1:function FindKeywords($t_{p},a^{\text{adv}}_{p},Q_{r}^{\text{adv}},{\bm{K}}$)
2:  $W\leftarrow$ All words in titles and abstracts from $Q_{r}^{\text{adv}}$.
@@
5:  keywordsSimilarity $\leftarrow$ $s((t_{p},a^{\text{adv}}_{p}),Q_{r}^{\text{adv}})$ $\triangleright$ Keeps track of estimated similarity to $r$
@@
8:      ${a^{\text{adv}}_{p}}^{(j)}\leftarrow$ adversarial abstract with words appended at the end $a^{\text{adv}}_{p}{}$} $w_{\max}^{(0)}$} … $w_{\max}^{(i-1)}$} $w_{j}$}”.
@@
10:   Let $w_{\max}^{(i)}$ be a word $w_{j}$ such that the associated $s((t_{p},{a^{\text{adv}}_{p}}^{(j)}),Q_{r}^{\text{adv}})$ is the highest in $W$.
11:   if $\underset{j}{\max}\ s((t_{p},{a^{\text{adv}}_{p}}^{(j)}),Q_{r}^{\text{adv}})<$ keywordsSimilarity then $\triangleright$ Stop if similarity does not increase
@@
15:   keywordsSimilarity $\leftarrow\underset{j}{\max}\ s((t_{p},{a^{\text{adv}}_{p}}^{(j)}),Q_{r}^{\text{adv}})$ $\triangleright$ Update estimated similarity with $w_{\max}^{(i)}$ added
@@
Appendix B Attack Budgets
We evaluate success rates under varying attack budgets by studying how the hyperparameters of IncludeThemes and InsertKeywords operations ($\bm{N}$, $\bm{M}$}, $\bm{K}$}) affect automatic attack success rates of 50 samples with natural rankings of 101 from the NeurIPS 2023 dataset.
First, Figure 6 shows attack success rates for varying $\bm{N}$—the number of IncludeThemes rounds— without using InsertKeywords (${\bm{M}}=0$, ${\bm{K}}=0$). Success rates generally increase with $\bm{N}$. This is expected because language model embeddings can be sensitive to paraphrasing, so taking the most-similar attempt amongst stochastic outputs helps improve attack success rates. However, there are signs of diminishing returns as $\bm{N}$ increases.
Second, in Figure 7, we report attack success rates for varying $\bm{M}$} and $\bm{K}$} in the InsertKeywords operation and FindKeywords subroutine, without using IncludeThemes ($\bm{N}=0$). The success rates increase with the total number of inserted keywords, ${\bm{M}}\times{\bm{K}}$, as shown by the color gradient from bottom left to top right in each subfigure. For a fixed keyword count, inserting smaller batches over more InsertKeywords iterations outperforms larger batches over fewer iterations, as seen in the contrast between upper-left and bottom-right squares in those subfigures. This supports the intuition that gradual updates help greedy search adapt to changes in $a^{\text{adv}}_{p}{}$ from earlier batches.
Figure 6: Attack success rates generally increase with $\bm{N}$ when the most similar attempt is kept out of $\bm{N}$ versions of $a^{\text{adv}}_{p}$ in IncludeThemes. The band is the standard error.
(a) Top-1
(b) Top-3
(c) Top-5
Appendix C Similarity Scores and Rankings
In this work, we evaluate attack success using natural and manipulated rankings, as relative similarity rankings are more broadly applicable across conferences that may use different optimization programs for reviewer assignment (Section 2.1). To relate natural rankings to absolute similarity, Figure 8 plots the mean similarity score (before attack) associated with each ranking across all papers in the NeurIPS 2023 dataset. For each paper, we rank all non-author reviewers by similarity and average the scores at each rank position over all papers.
Figure 8(a) shows that only a small fraction of reviewers have exceptionally high or low similarity scores to each paper, while mid-ranking reviewers exhibit a much more gradual decline in similarity scores. Figure 8(b) zooms in on selected rankings between 1 and 1001 for closer inspection.
(a) Mean similarity scores associated with each ranking from 1 to 7900. The shaded area represents standard error but is too small to be visible.
(b) Zoomed in plot of the mean similarity scores, only for selected rankings between 1 and 1001. Error bars represent standard errors.
Figure 8: Mean similarity scores (before attack) associated with different rankings across all papers in the NeurIPS 2023 dataset. The curve exhibits a steep decline at the highest and lowest ranks, indicating that only a small fraction of reviewers has exceptionally high or low similarity to a given paper.
Appendix D Similarity Scores and Attack Results
Some papers–such as those in niche fields–may have fewer high‐similarity reviewers, while others may have more. As a result, even colluding reviewer-paper pairs with identical ranking positions may have different values of similarity scores, leading to varying attack success rates. In this section, we explore how those natural similarity scores affect the likelihood of a successful attack.
For each paper, we calculate the similarities to all reviewers and then rank them. This allows us to compute the top and bottom quartiles with respect to the similarity scores of all reviewer-paper pairs with the same natural ranking. In particular, we focus on rankings 20, 101, 501, and 1001. We then revisit the main results from Section 5.3, applying this new stratification. For each ranking, we consider the colluding pairs identified in Section 5.3 and categorize them into two groups: those whose natural similarity falls in the bottom quartile and those in the top quartile. In Table 7, success rates are generally lower for the bottom quartile group, though attacks can still achieve considerable success in these cases. We note that results are mixed when the natural ranking is 20, likely because attacks are highly successful for both groups.
Table 7: Attack success rates of groups of colluding pairs with natural similarity (before attack) in the top and bottom quartiles for their respective natural rankings. The baseline attack success rates from Section 5.3 are also provided.
Attack Success Rates ($\pm$ SE)
Natural Top-K From Bottom quartile Top quartile
Ranking Section 5.3
[…]
Appendix E LLM Prompts
E.1 IncludeThemes Prompts
@@
E.1.1 Human-in-the-loop Mode
@@ E.1.2 Automatic Mode
For both human-in-the-loop and manual modes, we provide two examples to the LLM to guide its generation. The examples can be found on Zenodo https://doi.org/10.5281/zenodo.15588237.
E.2 The InsertKeywords Prompt
@@
For this operation, we also provide two examples to the LLM to guide its generation. The examples can be found on Zenodo https://doi.org/10.5281/zenodo.15588237.
E.3 Prompts used in Section 5.7 (Human Subject Experiment)
As mentioned in Section 5.7, the prompts used for automatic abstract modification in the human subject experiment is tuned separately on the NeurIPS 2023 test data. Although in different words, the IncludeThemes prompt here has the same rules as the automatic mode prompt presented in Appendix E.1.2 about the number of sentences to add, the inspired by format to follow, and using a scientific writing style. Similarly, for the InsertKeywords prompt, the rules here and in Appendix E.2 about how to insert the keywords, use professional writing styles, and not to add completely unrelated keywords are the same.
IncludeThemes Prompt (human subject experiment, fully automatic mode): Please help me edit my abstract’s introduction to explain in one sentence that my work is inspired by the list of previous works in the provided JSON dictionary under the key ”related previous works”. Do not include the titles of previous works. Use the usual writing style of technical academic abstracts, avoid exaggerations and figurative language. Do not use flowery words or phrases such as “prowess”. In addition, please remove any identifiable information (e.g. GitHub URLs) in my abstract by simply replacing them with [omitted for de-identification]. I will provide you with a JSON dictionary with the following structure: { ”title”: my paper’s title, ”abstract”: my paper’s abstract ”related previous works”: [ ”title”: title1, ”abstract”: abstract1, …. ] } Format your answer into JSON with the following schema: { ”title”: title string (should be the same as original), ”abstract”: edited abstract with short introduction, explaining my work is inspired by the previous works provided }